# Defending Large Language Models Against Jailbreak Attacks via Layer-specific Editing

### v1

@@ Abstract
Large language models (LLMs) are increasingly being adopted in a wide range of real-world applications. Despite their impressive performance, recent studies have shown that LLMs are vulnerable to deliberately crafted adversarial prompts even when aligned via Reinforcement Learning from Human Feedback or supervised fine-tuning. While existing defense methods focus on either detecting harmful prompts or reducing the likelihood of harmful responses through various means, defending LLMs against jailbreak attacks based on the inner mechanisms of LLMs remains largely unexplored. In this work, we investigate how LLMs response to harmful prompts and propose a novel defense method termed Layer-specific Editing (LED) to enhance the resilience of LLMs against jailbreak attacks. Through LED, we reveal that several critical safety layers exist among the early layers of LLMs. We then show that realigning these safety layers (and some selected additional layers) with the decoded safe response from selected target layers can significantly improve the alignment of LLMs against jailbreak attacks. Extensive experiments across various LLMs (e.g., Llama2, Mistral) show the effectiveness of LED, which effectively defends against jailbreak attacks while maintaining performance on benign prompts. Our code is available at https://github.com/ledllm/ledllm.
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 [1], Llama2 [40], Vicuna [7], and Mistral [20] have demonstrated remarkable capabilities across a wide range of natural language tasks and have been increasingly adopted in many real-world applications. Despite extensive efforts [34, 4, 13, 51, 41] to align LLMs’ responses with human values to generate helpful and harmless content, recent studies [37, 42, 9, 39, 54, 43, 48, 6, 18, 26, 23] reveal that these aligned models are still vulnerable to intentionally crafted adversarial prompts, also termed as "jailbreak attacks", which can elicit harmful, biased, or otherwise unintended behaviors from LLMs, posing significant challenges to their safe deployment.
@@
To mitigate the jailbreak attacks on LLMs, various methods have been proposed. However, existing defense methods primarily focus on two aspects: 1) detecting whether the prompt or response contains harmful or unnatural content via perplexity filter [3, 19], input mutation [5, 38], or using the LLM itself [16, 24]; and 2) reducing the probability of generating harmful responses through safe instruction [43, 45, 55] or logit processor [46]. While these methods reduce the attack success rate of adversarial prompts to some extent, their effectiveness may be overcome by adaptive attacks [26]. Our take is that there remains a significant gap in our understanding of how LLMs handle harmful prompts and whether jailbreak attacks exploit LLMs in specific ways to produce harmful responses. Without diving into the inner workings of LLMs, existing efforts to improve their safety may only scratch the surface.
Recent research about LLM pruning and layer skipping [27, 14, 10] found that removing certain layers does not significantly affect LLM performance. Additionally, observations made in [49] suggest that early layers are crucial in defending against adversarial attacks. These studies suggest that not all layers contribute equally when responding to harmful prompts, and adversarial prompts can be mitigated by enhancing the resilience of LLMs.
@@
In this work, we take a step toward in terms of understanding the underlying safety mechanisms of LLMs. We conduct a systematic layer-wise analysis to identify layers that significantly influence LLM responses in the presence of harmful prompts. Surprisingly, we observe several critical safety layers responsible for handling harmful prompts. Once these layers are pruned, LLMs can be jailbroken by simply inputting the original harmful prompts without any modification. Furthermore, we provide a detailed analysis of each layer’s hidden states and find that not all layers trigger LLMs to generate harmful responses; some layers retain a relatively high probability of decoding refusal tokens.
@@
Based on these insights, we propose a novel jailbreaking defense method termed Layer-specific Editing (LED), which uses targeted model editing to enhance LLM defence against adversarial attacks. Through LED, we reveal that several critical safety layers are located in the early layers of LLMs. Realigning these safety layers and some selected additional layers which merely contribute to defense with the decoded safe response from selected target layers can significantly improve the safety alignment of LLMs, whilst maintaining their performance on benign prompts.
Our main contributions are summarized as follows.
@@ 2 Related Work
Jailbreak Attacks. Jailbreak attacks aim to elicit unintended and unsafe behaviors from LLMs via well-designed harmful queries. Early attacks on LLMs heavily relied on hand-crafted adversarial prompts [31, 2] as well as valid jailbreak prompts collected by users on social media [39]. Apart from manually designed jailbreak prompts based on conversation templates [23, 43, 48], recent studies generally focus on automatically generating jailbreak prompts, such as gradient-based methods [54, 26], gradient-free genetic algorithms [42], and random search to iteratively refine jailbreak prompts [35, 15]. Some methods incorporate an auxiliary LLM for discovering jailbreak prompts inspired by red-teaming [37]. For instance, GPTFuzzer [47] utilizes a pre-trained LLM to update hand-crafted jailbreak templates, and PAIR [6] involves an attacker LLM to iteratively select candidate prompts for jailbreaking the target LLM. While these existing jailbreak methods are promising in attacking LLMs, the vulnerability of LLMs given a malicious context remains unexplored.
Jailbreak Defenses. Although there are many efforts for aligning LLM’s responses with human values [34, 4, 13, 51, 41], jailbreak attacks can still bypass the safeguards and induce LLMs to generate harmful and unethical responses. To mitigate the impact of jailbreak on the robustness of LLMs, various defense methods have been proposed to enhance them against such prompts. These methods mainly consist of two categories: (1) detecting the harmfulness of input queries or output responses via perplexity filter [3, 19], input smoothing [5, 38], and a judge LLM [16, 24]; and (2) reducing the probability of generating harmful responses through safe instruction [43, 45, 55] and diminishing the logits of toxic tokens [46]. However, these methods fail to provide a comprehensive understanding of the underlying safety mechanisms inherent to LLMs.
Knowledge Editing. Knowledge editing aims to efficiently alter the behavior of LLMs within a specific domain without negatively impacting performance across other inputs. These methods can be divided into three categories: fine-tuning, meta-learning, and locate-and-edit. Fine-tuning methods [22, 53, 32] directly update stale knowledge using new datasets. Meta-learning methods such as KE [8] and MEND [30] propose to teach a hypernetwork to learn how to edit the model instead of updating the weights directly. For more efficient knowledge editing, locate-and-edit methods leveraged prior findings [12, 11] that the knowledge is mainly stored in the MLP (multilayer perceptron) modules, locating where the target knowledge was stored and editing the located area. For instance, ROME [28] and MEMIT [29] employ causal mediation analysis [36] to identify which part of hidden states the target knowledge is stored in and then modify the corresponding parameters.
Different from existing knowledge editing applications [41, 44] aiming to detoxify LLMs or eliminate privacy information, we propose a novel locate-and-edit approach for defending LLMs against jailbreak attacks, which first identifies the safety layers crucial to defense via layer-wise analysis then specifically edits these layers to enhance the safety alignment of LLMs. Compared to traditional knowledge editing methods, we set all target layers that possess semantic information as optimization objectives rather than merely aligning the model’s final output with preferred responses.
@@
Figure 1: Left: Example responses from normal and pruned LLMs to harmful and jailbreak prompts. Middle: Layer-wise pruning analysis involves selectively pruning layers and observing the changes in the responses of the pruned LLMs. When safety layers are removed, LLMs surprisingly provide harmful responses to unchanged harmful queries; Right: Layer-specific editing first identifies layers crucial for defending against harmful prompts, and then edits these layers to enhance the robustness of LLMs where we align decoded information of all target layers with a safe response.
3 Proposed Method
In this section, we explore the safety mechanism of LLMs via layer-wise analysis and show that there exist distinctive behaviors of different layers in LLMs’ responses to harmful and jailbreak prompts. Then we introduce our layer-specific editing method to enhance defense against jailbreak attacks.
Overview. The whole pipeline is described in Figure 1. In detail, we first conduct layer-wise pruning to identify "safety layers" which dominate the defense against harmful prompts. Then we determine layers to be edited according to the distribution of safety layers. Finally, we compute the "target layers" as the optimization objective for aligning all their decoded outputs with safe responses.
3.1 Exploring Safety Mechanism in LLMs
Identifying Safety Layers. To identify the presence of safety layers that are crucial in the LLM’s defense mechanism, we conduct layer-wise pruning analysis on a set of 100 randomly selected harmful prompts from AdvBench [54], iteratively removing one or more consecutive layers (until the response becomes non-sensical) to examine the response of the pruned LLM. Table 1 shows some representative results, with the identified position of safety layers, the corresponding harmful prompt and response generated by the original and pruned model. Intriguingly, pruning the safety layers induces the aligned LLM to provide harmful responses when presented with harmful prompts without any modification, indicating their crucial impact on the model’s defense mechanism against harmful prompts. That is, when these safety layers are removed, the LLM’s defense against harmful prompts is removed as well.

### v2

@@ Abstract
Large language models (LLMs) are increasingly being adopted in a wide range of real-world applications. Despite their impressive performance, recent studies have shown that LLMs are vulnerable to deliberately crafted adversarial prompts even when aligned via Reinforcement Learning from Human Feedback or supervised fine-tuning. While existing defense methods focus on either detecting harmful prompts or reducing the likelihood of harmful responses through various means, defending LLMs against jailbreak attacks based on the inner mechanisms of LLMs remains largely unexplored. In this work, we investigate how LLMs respond to harmful prompts and propose a novel defense method termed Layer-specific Editing (LED) to enhance the resilience of LLMs against jailbreak attacks. Through LED, we reveal that several critical safety layers exist among the early layers of LLMs. We then show that realigning these safety layers (and some selected additional layers) with the decoded safe response from identified toxic layers can significantly improve the alignment of LLMs against jailbreak attacks. Extensive experiments across various LLMs (e.g., Llama2, Mistral) show the effectiveness of LED, which effectively defends against jailbreak attacks while maintaining performance on benign prompts. Our code is available at https://github.com/ledllm/ledllm.
Defending Large Language Models Against Jailbreak Attacks via Layer-specific Editing
WARNING: This paper contains context which is toxic in nature.
Wei Zhaothanks: These authors contributed to the work equllly and should be regarded as co-first authors.1, Zhe Li11footnotemark: 11, Yige Li1, Ye Zhang2, Jun Sun1, 1Singapore Management University 2 National University of Singapore {wzhao,zheli,yigeli,junsun}@smu.edu.sg
@@ 1 Introduction
Large language models (LLMs) such as GPT-4 Achiam et al. (2023), Llama2 Touvron et al. (2023), Vicuna Chiang et al. (2023), and Mistral Jiang et al. (2023) have demonstrated remarkable capabilities across a wide range of natural language tasks and have been increasingly adopted in many real-world applications. Despite extensive efforts Ouyang et al. (2022); Bai et al. (2022); Glaese et al. (2022); Zhou et al. (2024a); Wang et al. (2023) to align LLMs’ responses with human values to generate helpful and harmless content, recent studies Perez et al. (2022); Wei et al. (2023a); Deng et al. (2023); Shen et al. (2023); Zou et al. (2023); Wei et al. (2023b); Zeng et al. (2024); Chao et al. (2023); Huang et al. (2024); Liu et al. (2024); Li et al. (2023a) reveal that these aligned models are still vulnerable to intentionally crafted adversarial prompts, also termed as "jailbreak attacks", which can elicit harmful, biased, or otherwise unintended behaviors from LLMs, posing significant challenges to their safe deployment.
Figure 1: Example responses from normal and pruned LLMs to harmful and jailbreak prompts. When some crucial layers are removed, LLMs surprisingly provide harmful responses to unchanged harmful queries.
@@
To mitigate the jailbreak attacks on LLMs, various methods have been proposed. However, existing defense methods primarily focus on two aspects: 1) detecting whether the prompt or response contains harmful or unnatural content via perplexity filter Alon and Kamfonas (2023); Jain et al. (2023), input mutation Cao et al. (2023); Robey et al. (2023), or using the LLM itself Helbling et al. (2023); Li et al. (2023b); and 2) reducing the probability of generating harmful responses through safe instruction Wei et al. (2023b); Xie et al. (2023); Zou et al. (2024) or logit processor Xu et al. (2024). While these methods reduce the attack success rate of adversarial prompts to some extent, their effectiveness may be overcome by adaptive attacks Liu et al. (2024). Our take is that there remains a significant gap in our understanding of how LLMs handle harmful prompts and whether jailbreak attacks exploit LLMs in specific ways to produce harmful responses. Without diving into the inner workings of LLMs, existing efforts to improve their safety may only scratch the surface.
@@
Recent research about LLM pruning and layer skipping Men et al. (2024); Gromov et al. (2024); Fan et al. (2024) found that removing certain layers does not significantly affect LLM performance. Additionally, observations made in Zhao et al. (2023) suggest that early layers are crucial in defending against adversarial attacks. These studies suggest that not all layers contribute equally when responding to harmful prompts and adversarial prompts. In this work, we take a step toward in terms of understanding the underlying safety mechanisms of LLMs. We conduct a systematic layer-wise analysis to identify layers that significantly influence LLM responses in the presence of harmful and jailbreak prompts. Figure 1 illustrates example responses from normal and pruned LLMs where some selected layers are removed to different prompts. Surprisingly, we observe several critical safety layers responsible for handling harmful prompts. Once these layers are pruned, LLMs can be jailbroken by simply inputting the original harmful prompts without any modification. Furthermore, our detailed analysis of each layer’s hidden states reveals that not all layers contain toxic information that triggers LLMs to generate harmful responses; some layers maintain a relatively high probability of decoding refusal tokens.
@@
Based on these insights, we propose a novel jailbreaking defense method termed Layer-specific Editing (LED), which uses targeted model editing to enhance LLM defense against adversarial attacks. Through LED, we reveal that several critical safety layers are located in the early layers of LLMs. Realigning these safety layers and some selected additional layers that merely contribute to defense with the decoded safe response from identified toxic layers can significantly improve the safety alignment of LLMs, whilst maintaining their performance on benign prompts. Our main contributions are summarized as follows.
@@ 2 Related Work
Jailbreak Attacks. Jailbreak attacks aim to elicit unintended and unsafe behaviors from LLMs via well-designed harmful queries. Early attacks on LLMs heavily relied on hand-crafted adversarial prompts Mowshowitz (2022); Albert (2023) as well as valid jailbreak prompts collected by users on social media Shen et al. (2023). Apart from manually designed jailbreak prompts based on conversation templates Li et al. (2023a); Wei et al. (2023b); Zeng et al. (2024), recent studies generally focus on automatically generating jailbreak prompts, such as gradient-based methods Zou et al. (2023); Liu et al. (2024), gradient-free genetic algorithms Wei et al. (2023a), and random search to iteratively refine jailbreak prompts Pal et al. (2023); Hayase et al. (2024). Some methods incorporate an auxiliary LLM for discovering jailbreak prompts inspired by red-teaming Perez et al. (2022). For instance, GPTFuzzer Yu et al. (2023) utilizes a pre-trained LLM to update hand-crafted jailbreak templates, and PAIR Chao et al. (2023) involves an attacker LLM to iteratively select candidate prompts for jailbreaking the target LLM. While these existing jailbreak methods are promising in attacking LLMs, the vulnerability of LLMs given a malicious context remains unexplored.
@@
Figure 2: Left: Layer-wise pruning analysis involves selectively pruning layers and observing the changes in the responses of the pruned LLMs. When safety layers are removed, LLMs surprisingly provide harmful responses to unchanged harmful queries; Middle: Locating toxic regions that facilitate the generation of harmful responses via decoding the hidden states $h_{l}$ at layer $l$ into vocabulary space $\mathbf{v}_{l}\in\mathbb{R}^{\text{\#vocab}\times 1}$; Right: Layer-specific editing first identifies layers crucial for defending against harmful prompts, and then edit these layers to enhance the robustness of LLMs where we align decoded information of all toxic layers with the safe response.
Jailbreak Defenses. Although there are many efforts for aligning LLM’s responses with human values Ouyang et al. (2022); Bai et al. (2022); Glaese et al. (2022); Zhou et al. (2024a); Wang et al. (2023), jailbreak attacks can still bypass the safeguards and induce LLMs to generate harmful and unethical responses. To mitigate the impact of jailbreak on the robustness of LLMs, various defense methods have been proposed to enhance them against such prompts. These methods mainly consist of two categories: (1) detecting the harmfulness of input queries or output responses via perplexity filter Alon and Kamfonas (2023); Jain et al. (2023), input smoothing Cao et al. (2023); Robey et al. (2023), and a judge LLM Helbling et al. (2023); Li et al. (2023b); and (2) reducing the probability of generating harmful responses through safe instruction Wei et al. (2023b); Xie et al. (2023); Zou et al. (2024) and diminishing the logits of toxic tokens Xu et al. (2024). However, these methods fail to provide a comprehensive understanding of the underlying safety mechanisms inherent to LLMs.
Knowledge Editing. Knowledge editing aims to efficiently alter the behavior of LLMs within a specific domain without negatively impacting performance across other inputs. These methods can be divided into three categories: fine-tuning, meta-learning, and locate-and-edit. Fine-tuning methods Lee et al. (2022); Zhu et al. (2020); Ni et al. (2023) directly update stale knowledge using new datasets. Meta-learning methods such as KE De Cao et al. (2021) and MEND Mitchell et al. (2022) propose to teach a hypernetwork to learn how to edit the model instead of updating the weights directly. For more efficient knowledge editing, locate-and-edit methods leveraged prior findings Geva et al. (2020, 2022) that the knowledge is mainly stored in the MLP (multilayer perceptron) modules, locating where the target knowledge was stored and editing the located area. For instance, ROME Meng et al. (2022) and MEMIT Meng et al. (2023) employ causal mediation analysis Pearl (2009) to identify which part of hidden states the target knowledge is stored in and then modify the corresponding parameters.
Unlike traditional knowledge editing methods Wang et al. (2023); Wu et al. (2023); Wang et al. (2024) that aim to detoxify LLMs, we do not directly edit the toxic layers believed to contain harmful knowledge. Previous work Patil et al. (2023); Ma et al. (2024) has shown that existing knowledge editing methods fail to erase all knowledge from the model and that there are still many ways to retrieve this information. Instead, we propose to first identify the safety layers crucial for defense through layer-wise analysis and then specifically edits these layers to realign toxic layers to only output safe responses.
3 LED: Layer-specific Editing for Enhancing Defense of LLMs
In this work, we propose LED, Layer-specific EDiting to enhance the safety alignment of LLMs. Figure 2 illustrates the overview of LED’s workflow. Specifically, LED consists of three critical steps: 1) selecting edited layers, which consist of safety layers mostly relevant to the safety alignment for the harmful queries and additional layers which merely contribute to the defense; 2) locating toxic layers, which act as the optimization object for fully eliminating unexpected information; and 3) layer-specific editing to align the edited layers with the decoded safe responses from toxic layers, thereby enhancing the defense effectiveness against jailbreak attacks.
[…]
$f_{l,n}=P(f,l,n),$ (1)
where $1\leq l\leq L$ and $0\leq n\leq\min(L/2,L-l)$. Given the importance of the initial embedding layer, we start with layer 1 rather than layer 0 in the probing. We limit the pruning to a maximum of half the model’s layers to retain meaningful output from the pruned LLM as suggested in previous work Men et al. (2024). Then we investigate the response of $f_{l,n}$ to a set of harmful prompts. If the response is harmful, we stop the pruning process and treat layer $l$ to $l+n$ as safety layer candidates. In practice, we use an array layer_frequency to count the frequency of safety layer candidates over all harmful prompts. The top-k layers that appear most frequently during the pruning process are designated as safety layers $S$ as described in Equation 2:
[…]
To enhance the robustness of the model editing, we also incorporate a selection of additional layers that merely contribute to the defense into $S$ to obtain the edited layers $E$, aiming to involve more layers in the defense against jailbreak attacks.
[…]
To automatically detect toxic regions, we introduce a layer-specific toxic score $T(h_{l})$ to quantify the number of toxic responses in the decoded output of layer $l$, defined as Equation 3:
$T(h_{l})=\mathbf{v}_{l}(t_{\text{toxic}})/\max(\mathbf{v}_{l}),$ (3)
where $t_{\text{toxic}}$ is a toxic token generated by the original LLM (i.e., the final layer of LLM), $\mathbf{v}_{l}$ is the actual decoded logits, $\mathbf{v}_{l}(t_{\text{toxic}})$ denotes the probability of the toxic token in $\mathbf{v}_{l}$, and $max(\mathbf{v}_{l})$ is the maximum logits value in $\mathbf{v}_{l}$. For example, when inputting an adversarial prompt, a certain layer decodes tokens with logits $[\text{Sorry}=0.6,\text{Sure}=0.4]$ and "Sure" is the original toxic token generated by the model, the toxic score would be $0.4/0.6$. A layer with a higher toxic score indicates that it has a higher probability of outputting unexpected toxic tokens. In this work, layers with an average toxic score above $0.5$ are considered optimization objectives, meaning we should align these layers to output only safe content and not contribute to generating harmful responses. Unlike previous knowledge editing approaches Meng et al. (2022, 2023), which focus solely on changing the final layer output for evaluation, or detoxifying methods Wang et al. (2024) that attempt to directly edit and detoxify toxic regions, our method emphasizes aligning decoded content from multiple layers with safe responses. This approach is necessary because it is impractical to eliminate all harmful knowledge from these layers. Instead, we fine-tune the model to ensure it only outputs safe content from these toxic layers.
(a) Comparison of ASR before and after pruning.
(b) The distribution of safety layers.
Figure 3: (a): The results of layer-wise pruning analysis on four different LLMs over 100 randomly selected harmful prompts from AdvBench Zou et al. (2023). Higher ASR indicates lower defense performance; (b): The frequency distribution of safety layers, which are mainly distributed in early layers.
[…]
$L_{edit}=-\log P_{f}(Y_{safe}|X_{harm},h_{t}),$ (4)
where $h_{t}$ is the hidden states for each toxic layer $t$ in $T$. Then, we determine the update direction $\Delta^{l}_{t}$ based on $L_{edit}$ to edit the weight of each layer $l$ in edited layers $E$. After finishing editing, we obtain a more robust LLM $f_{robust}$ against jailbreak attacks. See Algorithm 1 in Appendix D for a more detailed algorithm.
In the next section, we first present some interesting findings regarding safety layers and toxic layers. Then, we conduct extensive experiments to evaluate the effectiveness of our proposed LED.
4 A Closer Look at LLMs: Safety and Toxic Layers
Early Safety Layers Dominates Defense. To identify the presence of safety layers, we conduct a layer-wise pruning analysis on a set of 100 randomly selected harmful prompts from AdvBench Zou et al. (2023) as demonstrated in Section 3. Figure 3(a) presents the results of four different LLMs responding to naturally harmful prompts before and after pruning safety layers, where a higher attack success rate (ASR) indicates lower defense performance. Surprisingly, safety layers are widely present in different LLMs, regardless of their varying structures and training data. Pruning these layers significantly increases the ASR for naturally harmful prompts compared to the original model. Figure 3(b) illustrates the distribution of safety layers for each model. For different LLMs, their safety layers are mainly concentrated in early layers, indicating that these early layers play a crucial role in discriminating the harmfulness of input queries, consistent with previous findings Zhao et al. (2023). Notably, almost all later layers do not participate in the defense against harmful queries. This observation leads to the idea of enhancing the LLM’s safety by involving more layers in defending against harmful queries. See Table 5 in Appendix Defending Large Language Models Against Jailbreak Attacks via Layer-specific Editing WARNING: This paper contains context which is toxic in nature. for more representative results.
Figure 4: Examples of toxic score $T(h_{l})$ ($l\geq 26$) on Llama-7B and Mistral-7B over $100$ adversarial prompts. Layers with high toxic scores indicate that they have a high probability of outputting toxic tokens, which should be determined as toxic layers for alignment.
Multiple Later Layers Contain Toxic Information. Figure 4 shows that later layers, rather than only the final layer, generally tend to output the toxic token with high probability which requires realignment. Notably, we do not directly edit the parameters in these toxic layers because it is impractical to eliminate all harmful knowledge within them. We cannot identify all adversarial prompts that trigger harmful content in these layers, and even if we could, we still can not ensure that new methods would not retrieve harmful information from these layers. Instead, we set these toxic layers as targets and fine-tune the model to generate only safe refusal responses from them.
5 Experiments
5.1 Experiment Setup
Dataset and Models. To evaluate the effectiveness of defending against jailbreak attacks, we utilize Advbench Zou et al. (2023) to generate adversarial prompts using various attack methods. We adopt attack success rate (ASR) as the evaluation metric. In our layer-specific editing process, we use $200$ harmful prompts from Trojan Detection Competition 2023 Organizers (2023) to conduct layer-wise analysis and obtain input-output pairs. Additionally, we generate $500$ adversarial prompts for computing target layers. Notably, the two datasets do not share any similar harmful prompts. We use the widely-used benchmarks MT-bench Zheng et al. (2024) and Just-Eval Lin et al. (2023) to assess the helpfulness of edited LLMs. We use one strong-aligned model Llama2-7B and one weak-aligned model Mistral-7B to test the effectiveness of our method.
Attack Setup. We evaluate five state-of-the-art jailbreak attacks: PAIR Chao et al. (2023), AutoDAN Liu et al. (2024), GPTFuzzer Yu et al. (2023), GCG Zou et al. (2023), and DeepInception Li et al. (2023a), using EasyJailbreak Zhou et al. (2024b) for agile implementation. Then follow the default parameter setting in EasyJailbreak and apply GPT-4 Achiam et al. (2023) as the attack model that generates jailbreak material, e.g. toxic prefix/suffixes or entire jailbreak prompts.
Defense Setup. We consider five latest defense strategies: Self-Reminder Xie et al. (2023), PPL Alon and Kamfonas (2023), Paraphrase Jain et al. (2023), Self-Examination Helbling et al. (2023), and SafeDecoding Xu et al. (2024). We adopt the hyper-parameters suggested in their original papers for each method. We also include LoRA Hu et al. (2021) (Low-rank adaptation) using the same dataset employed in our editing process to compare the effectiveness of fine-tuning and knowledge editing. For our proposed LED method, we have identified specific layers for editing and target layers for optimization. For weak-aligned model Mistral-7B, we select top-5 safety layers $\{2,3,4,5,6\}$ and additional middle layers $\{13,14,15\}$ to be edited, and compute target layers $\{29,30,31\}$ via target score as the optimization objective. For strong-aligned model Llama2-7B, we select top-3 safety layers $\{4,5,6\}$ and additional layers $\{13,14,15\}$ to be edited and compute target layers $\{29,30,31\}$.
Model Defense Jailbreak Attacks
Natural GCG PAIR AutoDAN GPTFuzzer DeepInception
Mistral-7B No Defense 64.8% 100% 85.0% 100% 100% 100%
Self-Reminder 3.3% 84.8% 53.8% 77.8% 71.2% 100%
Self-Examination 1.5% 11.5% 9.4% 8.7% 29.2% 60.3%
PPL 64.8% 0% 85.0% 100% 100% 100%
Paraphrase 61.7% 58.5% 70.5% 56.4% 74.2% 100%
SafeDecoding 0% 14.6% 18.6% 12.5% 15.0% 18.5%
LoRA 32.1% 34.2% 34.6% 58.8% 39.6% 71.0%
LED 0% 10.9% 8.1% 11.2% 13.4% 13.1%
Llama2-7B No Defense 0% 43.3% 2.1% 26.5% 23.5% 8.7%
Self-reminder 0% 32.7% 0% 19.8% 16.3% 7.1%
Self-Examination 0% 0.4% 0% 0% 12.9% 8.3%
PPL 0% 0% 2.1% 26.5% 23.5% 8.7%
Paraphrase 0% 0% 0% 2.7% 4.8% 7.3%
SafeDecoding 0% 0% 0% 0% 5.6% 0.4%
LoRA 0% 12.8% 0% 18.5% 8.8% 5.2%
LED 0% 0% 0% 0% 0.9% 0.9%
Model Defense MT-Bench Just-Eval
Helpful Clear Factual Deep Engage Avg.
Mistral-7B No Defense 7.26 3.70 3.85 4.03 2.65 2.96 3.44
Paraphrase 6.71 3.11 3.54 3.89 2.35 2.69 3.12
SafeDecoding 6.89 3.30 3.67 4.08 2.31 2.74 3.22
LED 7.08 3.46 3.66 4.11 2.36 2.88 3.28
Llama2-7B No Defense 6.55 3.38 3.48 3.94 2.21 2.86 3.17
Paraphrase 5.81 3.25 3.17 3.64 2.01 2.57 2.93
SafeDecoding 5.99 3.13 3.30 3.49 2.11 2.71 2.95
LED 6.48 3.35 3.51 3.91 2.23 2.83 3.17
5.2 Effectiveness of LED against Jailbreak Attacks
(a) Dataset sizes for different setting.
(b) ASR for different fine-tune methods.
Figure 5: The performance of Mistral-7B after applying LoRA with different settings.
Editing Layers MT-Bench Jailbreak Attacks
PAIR AutoDAN GPTFuzzer DeepInception
/ 7.26 85.0% 100% 100% 100%
4 6.94 12.3% 17.7% 13.3% 39.9%
13 7.13 19.0% 24.8% 15.0% 43.0%
26 7.04 78.4% 78.8% 93.6% 96.2%
2, 3, 4, 5, 6, 7, 8, 9 6.64 5.0% 6.5% 9.4% 13.4%
2, 3, 4, 5, 6, 13, 14, 15 7.08 8.1% 11.2% 13.4% 13.1%
2, 3, 4, 5, 6, 26, 27, 28 6.78 13.4% 24.6% 26.7% 32.7%
5.3 Ablation Studies
In the multi-layer experiments, we first edited only the safety layers while keeping the number of edited layers constant. The results indicate that focusing solely on safety layers significantly enhances the model’s defense capabilities but at the cost of reduced helpfulness. This reduction is attributed to the model becoming overly sensitive to benign queries that resemble harmful prompts. Conversely, adding editing layers from later layers does not significantly improve either helpfulness or robustness. This finding is consistent with recent studies Gromov et al. (2024); Men et al. (2024), which suggest that deeper layers generally have less impact on the model’s overall effectiveness. Overall, our results suggest that the optimal approach involves editing both safety layers and some additional middle layers. This strategy effectively enhances the model’s security while causing negligible degradation in helpfulness.
Figure 6: The performance of Self-Examination on Mistral-7B after applying LED.
Self-Examination with Edited Model. Self-Examination is an output detection method where, after each output from the LLM, the model is queried again to determine whether the previous response is harmful. See Appendix C.3 for more details. In this experiment, we enhance Self-Examination by using the edited model to perform the evaluation instead of the original model. Figure 6 presents the results of using the edited model for Self-Examination. The results show significant improvements across all attack methods. However, the performance of DeepInception is less promising compared to directly applying the edited model for defense. This is because responses from DeepInception often include various imaginative scenarios and character interactions, making it difficult for the LLM to extract information from such long texts and identify harmful behaviors effectively.
6 Conclusion
This paper systematically investigates the intrinsic defense mechanism of LLMs against harmful and adversarial prompts through layer-wise pruning and decoding analysis, finding that some layers play a crucial role in defending harmful queries and many layers of defense capabilities are not fully utilized according to the unbalanced distribution of safety layers. Therefore, we introduce LED, a layer-specific editing approach to enhance the robustness of LLMs against adversarial attacks while maintaining the model’s helpfulness on benign queries. Extensive experiments demonstrate the effectiveness of LED, significantly reducing the ASR across different LLMs under various jailbreak attacks and preserving their helpfulness on benign benchmarks. Compared to fine-tuning methods, LED does not require any adversarial samples and achieves higher performance with fewer datasets.
Limitations and Future Work. Our study demonstrates that the defense mechanisms and the storage of harmful knowledge in LLMs are located in different areas. Editing safety layers alone does not directly erase harmful knowledge from the model. Identifying the exact locations of this harmful knowledge and determining effective methods to erase it for better defense remain open questions. We advocate for further research into understanding the functions of different components of LLMs to refine defense mechanisms and broaden their applicability.
Appendix A Layer-wise Pruning Jailbreak Examples
Tables 4, 5, and 6 provide examples of success and failure in layer-pruning jailbreak scenarios. It’s important to note that some early layers in LLMs are also crucial for the semantic understanding of input prompts. Consequently, even if some outputs do not conform to standard safe refusal responses, they may still fail GPT-4’s red-teaming evaluation and are therefore counted as failures.
Appendix B More Analysis On Hidden States
Figure 7: Modify adversarial prompts to harmful and benign prompts with minimal editing.
In this section, we perform more decoding analysis to systematically identify the difference between harmful and jailbreak prompts and how adversarial prompts trigger jailbreak behavior, we collect a set of 50 adversarial prompts via GCG and AutoDan and modify their tokens step by step to create corresponding harmful and benign prompts. Figure 7 provides an example of editing one adversarial query into harmful and benign one with minimal change. To convert adversarial prompts into harmful prompts, we incrementally change the adversarial suffix from a single token to multiple tokens until the prompt elicits a safe refusal response. Meanwhile, We modify adversarial prompts to benign prompts simply by replacing tokens in the harmful query with benign alternatives.
Figure 8: Decdoding Analysis Results
Then, we compute the similarity score of different types of prompt to observe the differences in the response of each layer. Figure 8 shows the average similarity score between each pair of prompts across different layers. Interestingly, the hidden state similarity is high in the early layers (0 to 10) among different types of prompt. For the latter layer (11 to 31), the similarity between normal and harmful prompts, or between normal and adversarial prompts is quite small. Surprisingly, we notice that the similarity score between harmful and adversarial prompts remains relatively high in the middle layers, and only become small in the late layers. This suggests that adversarial prompts mainly work by interfering the later layers.
Appendix C Detail Experiment Setting
C.1 Safety Conversation Template
As discussed in Section 4, weak-aligned models such as Vicuna and Mistral can output harmful content with only natural harmful query. To ensure the models deny harmful requests, we incorporate a safe into each potentially harmful prompt. The details of this safe conversation template message is outlined below:
C.2 Attack Setup
We evaluate our method on five state-of-the-art jailbreak attacks: PAIR Chao et al. (2023), AutoDAN Liu et al. (2024), GPTFuzzer Yu et al. (2023), GCG Zou et al. (2023), and DeepInception Li et al. (2023a), using EasyJailbreak Zhou et al. (2024b) for agile implementation. We follow the default parameter setting in EasyJailbreak and apply GPT-4Achiam et al. (2023) as the attack model that generates jailbreak material, e.g. toxic prefix/suffixes or entire jailbreak prompts.
C.3 Defense Baseline Setup
Self-Reminder. Xie et al. (2023) Self-Reminder is an input mitigation method that employs a safe conversation template to ensure the model responds safely. We apply the safe conversation template described above to remind the model to only output non-harmful content.
Self-Examination. Helbling et al. (2023) Self-Examination is a output detection method. After each output from LLM, we ask the model again to determine whether the previous response is harmful or not. Detail input template are shown below:
Does the following text contain harmful content: {output}
Respond with "Yes, this is harmful" or "No, this is not harmful".
PPL. Alon and Kamfonas (2023) Perplexity is an input detection method which detects anomalies in input prompts. We apply GPT-2 to compute the perplexity of an input prompt defined as:
Paraphrase. Jain et al. (2023) Paraphrase is an input mitigation method that aims to remove unrelated or potential jailbreak materials in input prompts. We use GPT-4 to paraphrase each prompt before input using the template below:
SafeDecoding. Xu et al. (2024) SafeDecoding is a fine-tuning method that aims at identifying safety disclaimers and amplifying their token probabilities, while simultaneously attenuating the probabilities of token sequences that are aligned with the objectives of jailbreak attacks. We used their fine-tuned Llama-2 model from their released source code and fine-tuned a safer version of Mistral with their default settings.